# SIC-POVMs and orders of real quadratic fields

### v1

@@ 4.2. Galois actions on algebraic SICs
There is a well-defined action of $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ on algebraic complex lines which maps them to other algebraic complex lines. Given $\sigma\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and ${\mathbf{v}}\in\overline{\mathbb{Q}}^{d}$ we send ${\mathbf{v}}\mathbb{C}\mapsto\sigma({\mathbf{v}})\mathbb{C}$. This action is well-defined on algebraic complex lines, since if $\lambda\in\overline{\mathbb{Q}}^{\times}$, then ${\mathbf{v}}\mathbb{C}=\lambda{\mathbf{v}}\mathbb{C}$ while
@@
Let $\sigma_{c}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ denote complex conjugation and $E$ an algebraic Galois extension of $\mathbb{Q}$. Let $C(\sigma_{c}\!\mid\!E)$ be the centralizer of $\sigma_{c}$ restricted to $\operatorname{Gal}(E/\mathbb{Q})$, that is,
@@
For a line-SIC $S$, we will show there is a Galois field $E$ (depending on $S$) for which $C(\sigma_{c}\!\mid\!E)$ is such a restricted set. However, neither the field $F^{\rm proj}\!\left(S\right)$ nor $F^{\rm eproj}\!\left(S\right)$ is known to always be a Galois extension of $\mathbb{Q}$, although there is no known counterexample for $d\neq 3$. Thus, we give a name to the Galois closure of $F^{\rm eproj}\!\left(S\right)$.
@@
${\mathbf{\Pi}}(S)$ =$=\left\{\tfrac{1}{d}\Pi_{1},\tfrac{1}{d}\Pi_{2},\ldots,\tfrac{1}{d}\Pi_{d^{2}}\right\}.$ (4.5)
After rescaling if needed, all the entries of the $\mathbf{v}_{i}$ lie in $F^{\rm vec}\!\left(S\right)$, hence they lie in $F^{\rm proj}\!\left(S\right)$, so they lie in $F^{\rm Gproj}\!\left(S\right)$. The entries of the $\Pi_{i}$ also lie in $F^{\rm proj}\!\left(S\right)$, hence $F^{\rm Gproj}\!\left(S\right)$. Also,
@@
is the Hermitian projection onto $\sigma(\mathbf{v}_{i})$, so
@@
so ${\mathbf{\Pi}}(\sigma(S))$ is a SIC. Finally, $F^{\rm Gproj}\!\left(\sigma(S)\right)\subseteq F^{\rm Gproj}\!\left(S\right)$, because the entries of $\Pi_{i}$ lie in $F^{\rm Gproj}\!\left(S\right)$, and $F^{\rm Gproj}\!\left(S\right)$ is a Galois field fixed by $\sigma$. It follows by symmetry that $F^{\rm Gproj}\!\left(S\right)=F^{\rm Gproj}\!\left(\sigma^{-1}(\sigma(S))\right)\subseteq F^{\rm Gproj}\!\left(\sigma(S)\right)$ and thus $F^{\rm Gproj}\!\left(\sigma(S)\right)=F^{\rm Gproj}\!\left(S\right)$. ∎
@@ 4.3. Galois multiplets and super-equivalence classes of Weyl–Heisenberg SICs
We now assemble $\operatorname{PEC}(d)$-orbits of Weyl–Heisenberg SICs in dimension $d$ into multiplets under the Galois action by the subgroup $C(\sigma_{c}\!\mid\!F^{\rm eproj}\!\left(S\right))$ of $\operatorname{Gal}(F^{\rm eproj}\!\left(S\right)/\mathbb{Q})$ that commutes with complex conjugation. The coarser equivalence classes obtained under the combined geometric and Galois actions (i.e., unions of the orbits in a multiplet) will called Galois super-equivalence classes.
The conjecture of Appleby, Flammia, McConnell, and Yard [10, Conj. 2] implies that many of the fields we have been studying are equivalent Galois extensions of $\mathbb{Q}$.
@@
Assume 1.1 holds. Then in every dimension $d\geq 4$, for each Weyl–Heisenberg line-SIC $S$, the field $F^{\rm vec}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$, and
@@
By assertion (2) of 1.1 $F^{\rm vec}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$. Consequently, $F^{\rm vec}\!\left(S\right)$ is preserved by complex conjugation, so $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ holds by 3.5(1).
By assertion (2) of 1.1 $F^{\rm vec}\!\left(S\right)$ always contains the ray class field $H_{d^{\prime}\infty_{1}\infty_{2}}$ of $\mathbb{Q}(\Delta_{d})$. This class field must contain the class field of $\mathbb{Q}$ with conductor $d^{\prime}\infty$, which is $\mathbb{Q}(\zeta_{d^{\prime}})=\mathbb{Q}(\xi_{d})$. Therefore, $F^{\rm vec}\!\left(S\right)$ contains $F^{\rm proj}\!\left(S\right)(\xi_{d})=F^{\rm eproj}\!\left(S\right)$. Since $F^{\rm vec}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)\subseteq F^{\rm eproj}\!\left(S\right)$, the equality of those three fields follows. So $F^{\rm eproj}\!\left(S\right)$ is Galois and thus also equal to $F^{\rm Gproj}\!\left(S\right)$. ∎
We extract the following conjecture, of which all but the claim about $F^{\rm trip}\!\left(S\right)$ being Galois is already a consequence of 1.1.
Conjecture 4.10.
Fix a positive integer $d\neq 3$. For any $d$-dimensional Weyl–Heisenberg line SIC $S$, its $d^{2}$ vectors are always projective-algebraic. The associated fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ (and thus also $F^{\rm eproj}\!\left(S\right)$) are Galois extensions of $\mathbb{Q}$.
The evidence for a Galois action on Weyl–Heisenberg SICs was extensively detailed by Appleby, Yadsan-Appleby, and Zauner [4] in 2013.
The group $\operatorname{Gal}(\overline{\mathbb{Q}})/\mathbb{Q})$ acts on the elements of Weyl–Heisenberg group $\operatorname{WH}(d)$ coordinatewise on its matrices. Any automorphism $\sigma:\overline{\mathbb{Q}}\to\overline{\mathbb{Q}}$, sending a matrix $H\mapsto\sigma(H)$, acts as a homomorphism $\sigma(H_{1}H_{2})=\sigma(H_{1})\sigma(H_{2})$. This Galois action fixes the group $\operatorname{WH}(d)$ as a set, with any automorphism permuting its elements, since every Galois conjugate $\sigma(H)$ of a group element $H$ is also an element of $\operatorname{WH}(d)$.
By 3.8(2), $F^{\rm eproj}\!\left(S\right)$ is an invariant of the $\operatorname{PEC}(d)$ orbit of the Weyl–Heisenberg SIC $S$. If we assume 4.10, then $F^{\rm eproj}\!\left(S\right)$ is preserved by the Galois action by the subgroup $C(\sigma_{c}\!\mid\!F^{\rm eproj}\!\left(S\right))$ of $\operatorname{Gal}(F^{\rm eproj}\!\left(S\right)/\mathbb{Q})$. In other words, the extended projection field is invariant under both the geometric and the Galois actions.
We make the following definition.
@@
Assume 4.10 holds. For $d\geq 4$ the Galois multiplet of a $\operatorname{PEC}(d)$-equivalence class $[S]$ of Weyl–Heisenberg line-SICs is
$[[S]]=\{[\sigma(S)]:\sigma\in C(\sigma_{c}\!\mid\!F^{\rm eproj}\!\left(S\right))\},$ (4.10)
that is, the set of Galois equivalent orbits. We also define the equivalence relation $\approx_{GPEC}$, Galois super-equivalence, by
$S_{1}\approx_{GPEC}S_{2}\iff[[S_{1}]]=[[S_{2}]],$ (4.11)
and term the equivalence class of a SIC $S$ under $\approx_{GPEC}$ (that is, the union of the elements of $[[S]]$) the Galois super-equivalance class of $S$.
Remark 4.12.
Assuming 1.1 holds, this definition coincides with that of “Galois multiplet” in [9].
@@
Assume 4.10 holds. Then, the set $\operatorname{WHSIC}(d)$ is divided into a finite number of Galois super-equivalence classes under the equivalence relation $\approx_{GPEC}$. Each Galois super-equivalence class is a finite union of (geometric) $\operatorname{PEC}(d)$-equivalence classes.
Galois multiplets are studied in Sections 7 and 8 using data and more refined conjectures. The multiplets in a fixed dimension often have different sizes.
5. A conjecture on counting geometric equivalence classes of Weyl–Heisenberg SICs
In this section, we formulate a conjecture relating the number of SICs in dimensions $d\geq 4$ to class numbers of certain real quadratic orders. We present numerical evidence.
@@ 5.1. Class groups and class monoids of orders
For a more in-depth discussion of the nuances of orders, see [33] and [35, Ch. I, Sec. 12].
@@
For the SIC problem, we are specifically interested in the case when $\Delta=\Delta_{d}=(d+1)(d-3)$ for some integer $d$. In this case, the class monoid $\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})$ may be understood through bijections with other sets of number-theoretic interest.
@@
    the set $\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$, where $\mathcal{Q}(\Delta)$ is the set of binary quadratic forms of discriminant $\Delta$, and ${A}=\left(\begin{smallmatrix}s&t\\ u&v\end{smallmatrix}\right)\in\mathbf{GL}_{2}(\mathbb{Z})$ acts on the right by “twisted precomposition” $Q^{{A}}(x,y)=(\det{{A}})Q(sx+ty,ux+vy)$; and
@@
The bijection between (0) and (1) holds for quadratic orders in general. Specicially, by [33, Prop. A.12 and Prop. A.4], $\operatorname{Clm}(\mathcal{O}_{\Delta})=\operatorname{Clm}^{\flat}(\mathcal{O}_{\Delta})\cong\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$.
The bijection between (0) and (2) also holds for quadratic orders in general (and with appropriate modifications, nondegenerate quadratic rings) and is known as twisted Gauss composition. This is proven as [47, Thm. 1.2]; for an introduction to Gauss composition, see [16, Ch. 5].
We now exhibit a map $\phi:\mathcal{Q}(\Delta)\to\mathbf{GL}_{2}(\mathbb{Z})$ that induces a bijection between (2) and (3). For a quadratic form $Q(x,y)=ax^{2}+bxy+cy^{2}$ of discriminant $\Delta=(d+1)(d-3)$, define $\phi(Q):=\left(\begin{smallmatrix}\frac{1}{2}(b+d-1)&-a\\ c&\frac{1}{2}(-b+d-1)\end{smallmatrix}\right)$. Note that $\det(\phi(Q))=\frac{1}{4}(-b^{2}+(d-1)^{2})+ac=\frac{1}{4}(-\Delta+(d-1)^{2})=1$ and $\operatorname{Tr}(\phi(Q))=d-1$. It is straightforward to check that $\phi(Q^{R})={R}^{-1}\phi(Q){R}$ for ${R}\in\mathbf{GL}_{2}(\mathbb{Z})$. Thus, $\phi$ defines a map $\overline{\phi}:\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\to\mathcal{C}_{d-1}$. In the other direction, for ${A}=\left(\begin{smallmatrix}s&t\\ u&v\end{smallmatrix}\right)\in SL_{2}(\mathbb{Z})$, define the quadratic form $(\psi({A}))(x,y):=-tx^{2}+(s-v)xy+uy^{2}$. It is straightforward to check that $\operatorname{disc}(\psi({A}))=\Delta$ and $\psi({R}^{-1}{A}{R})=\psi({A})^{R}$, so $\psi$ defines a map $\overline{\psi}:\mathcal{C}_{d-1}\to\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$. A direct calculation shows $\overline{\phi}\circ\overline{\psi}=\operatorname{id}$ and $\overline{\psi}\circ\overline{\phi}=\operatorname{id}$. ∎

### v3

@@ 4.2. Galois actions on algebraic SICs
There is a well-defined action of $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ on algebraic complex lines which maps them to other algebraic complex lines. Given $\sigma\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and ${\mathbf{v}}\in\overline{\mathbb{Q}}^{d}$, we send ${\mathbf{v}}\mathbb{C}\mapsto\sigma({\mathbf{v}})\mathbb{C}$. This action is well-defined on algebraic complex lines, since if $\lambda\in\overline{\mathbb{Q}}^{\times}$, then ${\mathbf{v}}\mathbb{C}=\lambda{\mathbf{v}}\mathbb{C}$ while
@@
Let $\sigma_{c}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ denote complex conjugation and $E$ an algebraic Galois extension of $\mathbb{Q}$. Let $C(\sigma_{c}\!\mid\!E)$ denote the centralizer of $\sigma_{c}$ restricted to $\operatorname{Gal}(E/\mathbb{Q})$, that is,
@@
For a algebraic line-SIC $S$, we will show there is a Galois field $E$ (depending on $S$) for which $C(\sigma_{c}\!\mid\!E)$ is such a restricted set. Since neither the field $F^{\rm proj}\!\left(S\right)$ nor $F^{\rm eproj}\!\left(S\right)$ is known to always be a Galois extension of $\mathbb{Q}$, we give a name to the Galois closure of $F^{\rm eproj}\!\left(S\right)$.
@@
${\mathbf{\Pi}}(S)$ =$=\left\{\tfrac{1}{d}\Pi_{1},\tfrac{1}{d}\Pi_{2},\ldots,\tfrac{1}{d}\Pi_{d^{2}}\right\},$ (4.5)
where $\mathbf{v}_{i}$ are unit vectors. By definition of $F^{\rm proj}\!\left(S\right)$, the entries of all the $\Pi_{i}$ lie in $F^{\rm proj}\!\left(S\right)$, so they lie in $F^{\rm Gproj}\!\left(S\right)$. Also, using the commutativity of $\sigma$ and $\sigma_{c}$,
@@
is the Hermitian projection onto $\sigma(\mathbf{v}_{i})$, hence
@@
certifying that ${\mathbf{\Pi}}(\sigma(S))$ is a SIC. Finally, $F^{\rm Gproj}\!\left(\sigma(S)\right)\subseteq F^{\rm Gproj}\!\left(S\right)$, because the entries of $\Pi_{i}$ lie in $F^{\rm Gproj}\!\left(S\right)$, and $F^{\rm Gproj}\!\left(S\right)$ is Galois over $\mathbb{Q}$, hence $\sigma(F^{\rm Gproj}\!\left(S\right)=F^{\rm Gproj}\!\left(S\right)$. It follows by symmetry that $F^{\rm Gproj}\!\left(S\right)=F^{\rm Gproj}\!\left(\sigma^{-1}(\sigma(S))\right)\subseteq F^{\rm Gproj}\!\left(\sigma(S)\right)$ and thus $F^{\rm Gproj}\!\left(\sigma(S)\right)=F^{\rm Gproj}\!\left(S\right)$. ∎
@@ 4.3. Galois multiplets and super-equivalence classes of Weyl–Heisenberg SICs
We now assemble $\operatorname{PEC}(d)$-orbits of Weyl–Heisenberg SICs in dimension $d$ into multiplets under the Galois action by the subgroup $C(\sigma_{c}\!\mid\!F^{\rm Gproj}\!\left(S\right))$ of $\operatorname{Gal}(F^{\rm Gproj}\!\left(S\right)/\mathbb{Q})$ that commutes with complex conjugation. The coarser equivalence classes obtained under the combined geometric and Galois actions (i.e., unions of the orbits in a multiplet) will be called Galois super-equivalence classes.
The conjecture of Appleby, Flammia, McConnell, and Yard [9, Conj. 2] implies that many of the fields we have been studying are equivalent Galois extensions of $\mathbb{Q}$.
@@
Assume 1.1 holds. Then in every dimension $d\geq 4$, for each Weyl–Heisenberg line-SIC $S$, the field $F^{\rm vec}\!\left(S\right)$ is a finite Galois extension of $\mathbb{Q}$, and
@@
By 1.1(2), $F^{\rm vec}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$. Consequently, $F^{\rm vec}\!\left(S\right)$ is preserved by complex conjugation, so $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ holds by 3.5(1).
Also by 1.1(2), $F^{\rm vec}\!\left(S\right)$ always contains the ray class field $H_{d^{\prime}\infty_{1}\infty_{2}}$ of $\mathbb{Q}(\Delta_{d})$. This class field must contain the class field of $\mathbb{Q}$ with conductor $d^{\prime}\infty$, which is $\mathbb{Q}(\zeta_{d^{\prime}})=\mathbb{Q}(\xi_{d})$. Therefore, $F^{\rm vec}\!\left(S\right)$ contains $F^{\rm proj}\!\left(S\right)(\xi_{d})=F^{\rm eproj}\!\left(S\right)$. Since $F^{\rm vec}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)\subseteq F^{\rm eproj}\!\left(S\right)$, equality of those three fields follows. So $F^{\rm eproj}\!\left(S\right)$ is Galois and hence equal to $F^{\rm Gproj}\!\left(S\right)$. ∎
The evidence for a Galois action on Weyl–Heisenberg SICs was extensively detailed by Appleby, Yadsan-Appleby, and Zauner [4] in 2013. The group $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the elements of Weyl–Heisenberg group $\operatorname{WH}(d)$ coordinatewise on its matrices. Any automorphism $\sigma:\overline{\mathbb{Q}}\to\overline{\mathbb{Q}}$, sending a matrix $H\mapsto\sigma(H)$, acts as a homomorphism $\sigma(H_{1}H_{2})=\sigma(H_{1})\sigma(H_{2})$. This Galois action fixes the group $\operatorname{WH}(d)$ as a set, with any automorphism permuting its elements, since every Galois conjugate $\sigma(H)$ of a group element $H$ is also an element of $\operatorname{WH}(d)$.
By 3.8(2), $F^{\rm eproj}\!\left(S\right)$ is an invariant of the $\operatorname{PEC}(d)$ orbit of the Weyl–Heisenberg SIC $S$, hence so is its Galois closure $F^{\rm Gproj}\!\left(S\right)$. Thus the field $F^{\rm Gproj}\!\left(S\right)$ is invariant under both the geometric and the Galois actions.
The following definition coincides with “Galois multiplet” as defined in [8].
@@
For $d\geq 4$, the Galois multiplet of a $\operatorname{PEC}(d)$-equivalence class $[S]$ of Weyl–Heisenberg line-SICs is
$[[S]]=\{[\sigma(S)]:\sigma\in C(\sigma_{c}\!\mid\!F^{\rm Gproj}\!\left(S\right))\}.$ (4.10)
We also define an associated equivalence relation $\approx_{GPEC}$, termed Galois super-equivalence, by
$[S_{1}]\approx_{GPEC}[S_{2}]\iff[[S_{1}]]=[[S_{2}]],$ (4.11)
and term the union of the elements $[S_{i}]$ of $[[S]]$ the Galois super-equivalance class of $S$.
@@
Assume 1.1 holds. Then for $d\geq 4$ the set $\operatorname{WHSIC}(d)$ is partitioned into a finite number of Galois super-equivalence classes under the equivalence relation $\approx_{GPEC}$. Each Galois super-equivalence class is a finite union of (geometric) $\operatorname{PEC}(d)$-equivalence classes.
Galois multiplets in a fixed dimension often have different sizes.
5. Conjectures on counting geometric equivalence classes and Galois multiplets of Weyl–Heisenberg SICs
In this section, we formulate two conjectures relating the number of SICs in dimensions $d\geq 4$ to class numbers of certain real quadratic orders. Both these conjectures are numerically testable, and they are consequences of 1.2.
@@ 5.1. Class groups and class monoids of orders
For a more in-depth discussion of the nuances of orders, see [34] and [37, Ch. I, Sec. 12].
@@
For the SIC problem, we specifically consider the case when $\Delta=\Delta_{d}=(d+1)(d-3)$ for some integer $d$. In this case, the class monoid $\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})$ may be understood through bijections with other sets of number-theoretic interest.
@@
    the set $\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$, where $\mathcal{Q}(\Delta)$ is the set of binary quadratic forms of discriminant $\Delta$, and ${A}=\left(\begin{smallmatrix}s&t\\ u&v\end{smallmatrix}\right)\in\mathbf{GL}_{2}(\mathbb{Z})$ acts on the right by $Q^{{A}}(x,y)=(\det{{A}})Q(sx+ty,ux+vy)$; and
@@
The bijection between (0) and (1) holds for quadratic orders in general. Specifically, by Halter-Koch [28, Sec. 5.5], one has, as sets, $\operatorname{Clm}(\mathcal{O}_{\Delta})=\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$.
The bijection between (0) and (2) also holds for quadratic orders in general (and with appropriate modifications, nondegenerate quadratic rings) and is known as twisted Gauss composition. This is proven as [47, Thm. 1.2]; for an introduction to Gauss composition, see [15, Ch. 5].
We now exhibit a bijection between (2) and (3) which specifically holds only for $\Delta=(d+1)(d-3)$. It is defined by a map $\phi:\mathcal{Q}(\Delta)\to\mathbf{GL}_{2}(\mathbb{Z})$ that induces the bijection. For a quadratic form $Q(x,y)=ax^{2}+bxy+cy^{2}$ of discriminant $\Delta=(d+1)(d-3)$, define $\phi(Q):=\left(\begin{smallmatrix}\frac{1}{2}(b+d-1)&-a\\ c&\frac{1}{2}(-b+d-1)\end{smallmatrix}\right)$. Note that
$\det(\phi(Q))=\frac{1}{4}(-b^{2}+(d-1)^{2})+ac=\frac{1}{4}(-\Delta+(d-1)^{2})=1$ (5.4)
and $\operatorname{Tr}(\phi(Q))=d-1$. It is straightforward to check that $\phi(Q^{R})={R}^{-1}\phi(Q){R}$ for ${R}\in\mathbf{GL}_{2}(\mathbb{Z})$. Thus, $\phi$ defines a map $\overline{\phi}:\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\to\mathcal{C}_{d-1}$. In the other direction, for ${A}=\left(\begin{smallmatrix}s&t\\ u&v\end{smallmatrix}\right)\in\mathbf{SL}_{2}(\mathbb{Z})$, define the quadratic form $(\psi({A}))(x,y):=-tx^{2}+(s-v)xy+uy^{2}$. It is straightforward to check that $\operatorname{disc}(\psi({A}))=\Delta$ and $\psi({R}^{-1}{A}{R})=\psi({A})^{R}$, so $\psi$ defines a map $\overline{\psi}:\mathcal{C}_{d-1}\to\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$. A direct calculation shows $\overline{\phi}\circ\overline{\psi}=\operatorname{id}$ and $\overline{\psi}\circ\overline{\phi}=\operatorname{id}$. ∎